divide 630 between x:y in the ratio 1/4:1/6?
Answer by RAY100(1637)

x/y = (1/4)/ (1/6) = 6/4= 3/2,,,,,,(or 1.5)
.
3z+2z = 630
.
5z = 630
.
z= 126
.
.
3z=3*126=378 =x
.
2z=2*126=252 = y
.
check,,,,,,x/y = 378/252= 1.5
.
378 +252 = 630,,,,ok
.